# I3-MRec: Invariant Learning with Information Bottleneck for Incomplete Modality Recommendation

### v1

@@ Abstract.
Multimodal recommender systems (MRS) improve recommendation performance by integrating diverse semantic information from multiple modalities. However, the assumption of the availability of all modalities rarely holds in practice due to missing images, incomplete descriptions, or inconsistent user content. These challenges significantly degrade the robustness and generalization capabilities of current models. To address these challenges, we introduce a novel method called I3-MRec, which uses Invariant learning with Information bottleneck principle for Incomplete Modality Recommendation. To achieve robust performance in missing modality scenarios, I3-MRec enforces two pivotal properties: (i) cross-modal preference invariance, which ensures consistent user preference modeling across varying modality environments, and (ii) compact yet effective modality representation, which filters out task-irrelevant modality information while maximally preserving essential features relevant to recommendation. By treating each modality as a distinct semantic environment, I3-MRec employs invariant risk minimization (IRM) to learn modality-specific item representations. In parallel, a missing-aware fusion module grounded in the Information Bottleneck (IB) principle extracts compact and effective item embeddings by suppressing modality noise and preserving core user preference signals. Extensive experiments conducted on three real-world datasets demonstrate that I3-MRec consistently outperforms existing state-of-the-art MRS methods across various modality-missing scenarios, highlighting its effectiveness and robustness in practical applications. The code and processed datasets are released at https://github.com/HuilinChenJN/I3-MRec.
@@ 1. Introduction
Multimodal data has experienced explosive growth across various online platforms, including Amazon  [^1] https://www.amazon.com, YouTube [^2] https://www.youtube.com, and TikTok [^3] https://www.tiktok.com. To mitigate information overload and help users discover relevant items, Multimodal Recommender Systems (MRS) (Liu et al., 2023; Zhou et al., 2023a; Wang et al., 2023b; Wei et al., 2023b; Yu et al., 2023; Lin et al., 2024; Xu et al., 2024; Lin et al., 2025; Li et al., 2024) leverage rich information from various modalities (e.g., item images, review texts) to model user preferences and generate more accurate recommendations. However, most existing MRS methods assume the availability of complete modality information to effectively extract recommendation-relevant signals (Bai et al., 2024; Malitesta et al., 2024a). This assumption often fails in real-world settings, where incomplete modality data is often encountered in real-world applications, such as information retrieval (Wang et al., 2018; Wu et al., 2020b; Malitesta et al., 2024b), computer vision (Wang et al., 2023a), and video-sharing platforms (Li et al., 2025). As illustrated in Fig. 1, many existing multimodal recommendation models suffer from performance degradation or even fail when confronted with missing modalities [^4] For fair comparison, results are reproduced using the official code of the referenced method on the same dataset. . This frequent occurrence poses a substantial challenge to the robustness and practical applicability of MRS in real-world deployments.
@@
Although these methods have demonstrated promising effectiveness, the core challenges associated with missing modalities remain inadequately addressed: How should an optimal recommender system perform when confronted with modality missingness? On the one hand, generative methods (Lin et al., 2023; Li et al., 2025) often require extensive and time-consuming pre-training for each modality. This requirement contradicts the missing modality assumption, as it presumes the availability of large-scale, fully annotated multimodal data. On the other hand, existing representation-based approaches frequently overlook the distinct characteristics of each modality. For example, MILK (Bai et al., 2024) adopts a modality-wise mixup strategy to simulate modality-missing environments. However, such operations can disrupt modality-specific semantic structures, ultimately making the model more vulnerable to performance degradation under real-world missing modality conditions.
@@
  • Cross-modal preference invariance: User preferences are inherently stable across different modality sources (Du et al., 2022). A robust model should therefore learn preference-oriented multimodal representations that remain invariant across modalities. This ensures that each individual modality representation can sufficiently support accurate preference prediction, thereby mitigating the negative impact of missing modalities.
  • Compact yet effective modality representation: The learned item modality representations should maximally preserve user preference information pertinent to recommendation tasks while minimizing dependence on the raw modality content. This enables the model to maintain strong recommendation performance even in the absence of certain modalities.
These two principles together promote recommendation robustness by ensuring robust representations and reducing the model’s dependency on potentially unavailable modality data. To address this critical challenge, we propose I3-MRec (Invariant learning with Information Bottleneck for Incomplete-Modality Recommendation), a unified framework specifically designed for robust recommendation under modality-missing scenarios. To achieve the goal, I3-MRec adopts an invariant learning framework that explicitly treats each modality as a distinct semantic environment. It leverages a GCN-based architecture combined with Invariant Risk Minimization (IRM) to learn robust user and item representations (thereby satisfying the first property). Furthermore, we introduce a missing-aware fusion module guided by the Information Bottleneck (IB) principle (Wei et al., 2022; Yang et al., 2024, 2025), which selectively retains preference-relevant modality information while discarding irrelevant noise (Du et al., 2022) (in alignment with the second property).
@@
  • We introduce a principled invariant learning approach that models each modality as a distinct semantic environment. This enables learning of modality-invariant and semantically consistent user and item representations, improving resilience to missing modality inputs.
  • We design a missing-aware fusion module guided by the information bottleneck principle, which selectively retains preference-relevant information while discarding modality noise, resulting in compact yet effective item representations.
@@ 2.1. Multimodal Recommendation and Modality Missingness
Multimodal Recommender Systems leverage various multimedia content to enhance user-item interactions, offering significant advantages in applications such as online shopping (Kang et al., 2017; Liu et al., 2022), video sharing platforms (Wei et al., 2019; Cai et al., 2022), and social networks (Zhou et al., 2023b). Early approaches (Chen et al., 2017; He and McAuley, 2016; Kang et al., 2017) incorporated multimodal data as auxiliary features to enhance item representations. For instance, VBPR (He and McAuley, 2016) integrates visual embeddings with item ID embeddings to better model user preferences. With the development of Graph Neural Networks (GNNs)(He et al., 2020; Wang et al., 2019; Liu et al., 2021, 2025), subsequent work extended these models to incorporate multimodal signals into graph-based learning(Wei et al., 2019, 2020; Ma et al., 2024; Cai et al., 2021), improving the quality of user and item node representations. Some methods, such as LATTICE (Zhang et al., 2021), capture item-item semantic affinities via modality-aware graphs, while DualGNN (Wang et al., 2021) models user-user relations to reveal latent preference patterns. In addition, recent advances have incorporated attention mechanisms (Liu et al., 2022), self-supervised learning (Tao et al., 2022; Zhou et al., 2023a), and contrastive learning (Lin et al., 2023; Wei et al., 2023a) to better align and integrate cross-modality information.
Despite their effectiveness, these methods typically assume the availability of complete modality information. In practice, missing modalities can lead to significant performance degradation (Malitesta et al., 2024a; Bai et al., 2024). To address this challenge, existing solutions fall into two broad categories: Generation-based approaches aim to reconstruct missing modality features through modality generation techniques (Wang et al., 2018; Wu et al., 2020b; Bai et al., 2024; Ganhör et al., 2024; Lin et al., 2023; Li et al., 2025). While effective in controlled settings, these methods typically incur high computational costs and lack flexibility when handling random or modality-specific missing patterns. e.g., MoDiCF (Li et al., 2025) requires an auxiliary diffusion model to reconstruct missing modalities. Representation Learning-based methods focus on directly learning robust representations that remain effective even when specific modalities are missing (Bai et al., 2024; Malitesta et al., 2024b). For instance, MILK (Bai et al., 2024) introduces a modality-invariant learning strategy; however, it assumes complete modality availability during training, which limits its applicability in real-world scenarios where missing modalities are inherent and unpredictable. By formulating the task within an invariant learning framework and incorporating an IB-guided fusion strategy, I3-MRec generalizes effectively across various modality-missing scenarios.
@@ 2.3. Information Bottleneck
The Information Bottleneck (IB) principle, based on information theory, is widely applied in machine learning tasks such as model robustness (Wu et al., 2020a; Zhang et al., 2022; Ding et al., 2023), fairness (Gronowski et al., 2023), explainability (Espinosa Zarlenga et al., 2022; Wang et al., 2023c), and recommendation systems (Wei et al., 2022; Yang et al., 2024, 2025). For input data $X$, hidden representation $Z$, and prediction label $Y$, IB principle suggests that an effective representation retains minimal sufficient information for the prediction task (Tishby et al., 2000; Tishby and Zaslavsky, 2015; Hu et al., 2024): $Max:I(Y;Z)-\beta I(X;Z)$. where $I(Y;Z)$ and $I(X;Z)$ are the mutual information between variables, with $\beta$ balancing the two terms. Calculating mutual information for continuous variables is difficult, especially in deep learning. Approximations using neural networks, such as MINE (Belghazi et al., 2018), InfoNCE (Oord et al., 2018), and variational methods (Tishby and Zaslavsky, 2015), are commonly used. Recently, CLUB (Cheng et al., 2020) has emerged to estimate the upper bound of mutual information using a log-ratio contrastive loss, which is more general for high-dimensional tasks without prior assumptions.
@@ 3. Preliminaries
Figure 2. Overview of the proposed I3-MRec framework. The model first learns modality-specific user and item representations using a graph-based approach, guided by IRM to ensure invariance across modalities. Then, an IB-based fusion module generates compact yet effective representations by maximizing preference-relevant information while discarding modality noise.
@@ 3.2. Modality Missingness Problem
MRS typically models user preferences by first extracting modality-specific features and then fusing them into a unified item representation. Formally, this process can be described as:
@@ 4.1.1. Collaborative Modality Graph Learning
where $\bm{A}$ is the adjacency matrix built from user-item interactions $\mathcal{R}$ and its transpose $\mathcal{R}^{T}$, and $\bm{D}$ is the corresponding diagonal degree matrix. By aggregating node features across $L$ propagation layers, the final modality-specific embeddings can be obtained:
@@
Here, $\bm{E}^{k}=\{\bm{E}^{k}_{U},\bm{E}^{k}_{I}\}$ represents modality-specific representations of users and items. Based on the learned representations, we can estimate the preference score of the user $u$ for item $i$ in the $k$-th modality environment:
@@ 4.1.2. Invariant Learning Optimization
After obtaining predictions across $K$ modality environments, we explicitly encourage the consistency and stability of the learned representations across modalities. Specifically, we employ Invariant Risk Minimization  (Arjovsky et al., 2019; Ahuja et al., 2020; Du et al., 2022) to guide the learning of modality-specific representations that remain invariant across environments. The corresponding optimization objective is formulated as follows:
@@ 4.2. Missing-aware Feature Fusion
To further obtain the robust item representation for predicting user preference, we design a missing-aware feature fusion module grounded in the Information Bottleneck (IB) principle, as illustrated in the lower part of Figure 3. This module is encouraged to preserve preference-relevant modality semantics while compressing noisy or redundant information, enabling the generation of compact and robust item representations resilient to missing modalities.
@@ 4.2.1. Modality-Missing Scenarios Simulation
This strategy exposes the proposed model to various modality-missing scenarios during training, thus encouraging the fusion module to learn robust item representations.
@@ 4.2.2. Information Bottleneck-guided Feature Fusion
Given generated representations, we employ the Information Bottleneck principle to guide the optimization of the feature fusion module. It explicitly balances mutual information maximization and minimization, encouraging the model to retain task-relevant signals while compressing modality-specific noise. The corresponding optimization objective is defined as:
@@
Term1: Maximizing Mutual Information with Recommendation Signals. This term aims to maximize the mutual information between item representations and the recommendation target. Given the user embedding $e_{u}$ and the generated item representation $z_{i}$, the predicted preference score is computed as $\bar{r}_{u,i}^{k}=e_{u}^{T},z_{i}$. We select popular BPR ranking loss  (Rendle et al., 2009) to optimize this term:
@@
Term3: Minimizing Mutual Information for Redundancy Reduction. This term aims to compress modality-specific information by minimizing the mutual information between the generated representation and the raw modality input. However, directly computing mutual information between two high-dimensional variables is intractable. To address this, we adopt the Contrastive Log-ratio Upper Bound (CLUB) (Cheng et al., 2020) in our work, which uses a variational distribution $q_{\phi}^{k}(\cdot|\cdot)$ for each modality to approximate the mutual information:
(14) $\mathcal{L}_{comp}=\frac{1}{|K|}\sum_{k\in\{a,t,v\}}{\sum_{i\in\mathcal{I}}\left[\log q_{\phi^{k}}^{k}(z_{i}|e_{i}^{k})-\log q_{\phi^{k}}(z_{i}|x_{i}^{k})\right]}.$
here, $q_{\phi}^{k}(\cdot|\cdot)$ is implemented as a two-layer MLP with modality-specific parameter $\phi^{k}$, optimized in a sample-based manner. Minimizing $\mathcal{L}^{comp}$ effectively reduces the task-irrelevant information, yielding a more compact and effective representation.
@@
This unified loss function integrates four components: the recommendation loss ($\mathcal{L}_{rec}$), the invariant representation learning loss ($\mathcal{L}_{IRM}$), the preference consistency loss ($\mathcal{L}_{pcon}$), and the information compression loss ($\mathcal{L}_{comp}$). The hyperparameters $\alpha$ and $\beta$ control the relative importance of the mutual information regularization terms. By jointly optimizing these objectives, the model learns robust and compact preference representations that generalize well to modality-missing scenarios.

### v2

@@ Abstract.
Multimodal recommender systems (MRS) improve recommendation performance by integrating complementary semantic information from multiple modalities. However, the assumption of complete multimodality rarely holds in practice due to missing images and incomplete descriptions, hindering model robustness and generalization. To address these challenges, we introduce a novel method called I3-MRec, which uses Invairant learning with Information bottleneck principle for Incomplete Modality Recommendation. To achieve robust performance in missing modality scenarios, I3-MRec enforces two pivotal properties: (i) cross-modal preference invariance, ensuring consistent user preference modeling across varying modality environments, and (ii) compact yet effective multimodal representation, as modality information becomes unreliable in such scenarios, reducing the dependence on modality-specific information is particularly important. By treating each modality as a distinct semantic environment, I3-MRec employs invariant risk minimization (IRM) to learn preference-oriented representations. In parallel, a missing-aware fusion module is developed to explicitly simulate modality-missing scenarios. Built upon the Information Bottleneck (IB) principle, the module aims to preserve essential user preference signals across these scenarios while effectively compressing modality-specific information. Extensive experiments conducted on three real-world datasets demonstrate that I3-MRec consistently outperforms existing state-of-the-art MRS methods across various modality-missing scenarios, highlighting its effectiveness and robustness in practical applications.
@@ 1. Introduction
Multimodal data has experienced explosive growth across various online platforms, including Amazon  [^1] https://www.amazon.com, YouTube [^2] https://www.youtube.com, and TikTok [^3] https://www.tiktok.com. To mitigate information overload and help users discover relevant items, Multimodal Recommender Systems (MRS) (Liu et al., 2023a; Zhou et al., 2023a; Wang et al., 2023b; Wei et al., 2023b; Yu et al., 2023; Lin et al., 2024; Xu et al., 2024; Lin et al., 2025; Li et al., 2024) leverage rich information from various modalities (e.g., item images, review texts) to model user preferences and generate more accurate recommendations. However, most existing MRS methods assume the availability of complete modality information to effectively extract recommendation-relevant signals (Bai et al., 2024; Malitesta et al., 2024a). This assumption often fails in real-world settings, where incomplete modality data is often encountered in real-world applications, such as information retrieval (Wang et al., 2018; Wu et al., 2020b; Malitesta et al., 2024b), computer vision (Wang et al., 2023a), and video-sharing platforms (Li et al., 2025). As illustrated in Fig. 1, many existing multimodal recommendation models suffer from performance degradation or even fail when confronted with missing modalities [^4] For fair comparison, results are reproduced using the official code of the referenced method on the same dataset. . This frequent occurrence poses a substantial challenge to the robustness and practical applicability of MRS in real-world deployments.
@@
Although these methods have demonstrated promising effectiveness, the core challenges associated with missing modalities remain inadequately addressed: How should an optimal recommender system perform when confronted with modality missingness? On the one hand, generative methods (Lin et al., 2023; Li et al., 2025) often require extensive and time-consuming pre-training for each modality. This requirement contradicts the missing modality assumption, as it presumes the availability of large-scale, fully annotated multimodal data. On the other hand, existing representation-based approaches overlook the distinct characteristics of each modality. For example, MILK (Bai et al., 2024) adopts a modality-wise mixup strategy to simulate modality-missing environments. However, such operations can disrupt modality-specific semantic structures, ultimately making the model more vulnerable to performance degradation under real-world missing modality conditions.
@@
  • Cross-modal Preference Invariance: User preferences are inherently stable across different modality sources (Du et al., 2022). A robust model should therefore learn preference-oriented user/item representations that remain invariant across modalities. This ensures that each individual modality representation can sufficiently support accurate preference prediction, thereby mitigating the negative impact of missing modalities.
  • Compact yet Effective Multimodal Representation: The learned multimodal representations should maximally preserve user preference information pertinent to recommendation tasks while minimizing dependence on the raw modality information. As modality information becomes unreliable in such scenarios, especially when missing modalities randomly occur during training, reducing the dependence on modality-specific information is particularly important (Ma et al., 2021; Woo et al., 2023; Saeed et al., 2024).
These two principles together promote recommendation robustness by ensuring robust representations and reducing the model’s dependency on potentially unavailable modality data. To address this critical challenge, we propose I3-MRec (Invariant learning with Information Bottleneck for Incomplete-Modality Recommendation), a unified framework specifically designed for robust recommendation under modality-missing scenarios. To achieve the goal, I3-MRec adopts an invariant learning framework that explicitly treats each modality as a distinct semantic environment. It leverages a GCN-based architecture combined with Invariant Risk Minimization (IRM) to learn robust user and item representations (thereby satisfying the first property). Furthermore, we introduce a missing-aware fusion module guided by the Information Bottleneck (IB) principle (Tishby et al., 2000; Tishby and Zaslavsky, 2015; Hu et al., 2024) to explicitly simulate diverse modality-missing scenarios during training, maximizing mutual information between multimodal representations across these scenarios while minimizing the mutual information between the multimodal representation and each raw-modality feature. (in alignment with the second property). We have released the code and relevant parameter settings to facilitate repeatability as well as further research [^5] https://github.com/HuilinChenJN/I3-MRec..
@@
  • We introduce a principled invariant learning approach that models each modality as a distinct semantic environment. This enables learning of preference-oriented user/item representations, improving resilience to missing modality inputs.
  • We design a missing-aware fusion module guided by the Information Bottleneck principle, which selectively retains preference-relevant information while compressing modality-specific information, resulting in compact yet effective multimodal representations.
@@ 2.1. Multimodal Recommendation and Modality Missingness
Multimodal Recommender Systems leverage various multimedia content to enhance user-item interactions, offering significant advantages in applications such as online shopping (Kang et al., 2017; Liu et al., 2022), video sharing platforms (Wei et al., 2019; Cai et al., 2022), and social networks (Zhou et al., 2023b). Early approaches (Chen et al., 2017; He and McAuley, 2016; Kang et al., 2017) incorporated multimodal data as auxiliary features to enhance item representations. For instance, VBPR (He and McAuley, 2016) integrates visual embeddings with item ID embeddings to better model user preferences. With the development of Graph Neural Networks (GNNs)(He et al., 2020; Wang et al., 2019; Liu et al., 2021, 2025), subsequent work extended these models to incorporate multimodal signals into graph-based learning(Wei et al., 2019, 2020; Ma et al., 2024; Cai et al., 2021), improving the quality of user and item node representations. Some methods, such as LATTICE (Zhang et al., 2021), capture item-item semantic affinities via modality-aware graphs. while DualGNN (Wang et al., 2021) models user-user relations to reveal latent preference patterns. In addition, recent advances have incorporated attention mechanisms (Liu et al., 2022), self-supervised learning (Tao et al., 2022; Zhou et al., 2023a), and contrastive learning (Lin et al., 2023; Wei et al., 2023a) to better align and integrate cross-modality information.
Despite their effectiveness, these methods typically assume the availability of complete modality information. In practice, missing modalities can lead to significant performance degradation (Malitesta et al., 2024a; Bai et al., 2024). To address this challenge, existing solutions fall into two broad categories: Generation-based approaches aim to reconstruct missing modality features through modality generation techniques (Wang et al., 2018; Wu et al., 2020b; Bai et al., 2024; Ganhör et al., 2024; Lin et al., 2023; Li et al., 2025). While effective in controlled settings, these methods typically incur high computational costs and lack flexibility when handling random or modality-specific missing patterns. e.g., MoDiCF (Li et al., 2025) requires an auxiliary diffusion model to reconstruct missing modalities. Representation Learning-based methods focus on directly learning robust representations that remain effective even when specific modalities are missing (Bai et al., 2024; Malitesta et al., 2024b). For instance, MILK (Bai et al., 2024) introduces a modality-invariant learning strategy; however, it assumes complete modality availability during training, which limits its applicability in real-world scenarios. By formulating the task within an invariant learning framework and incorporating an IB-guided fusion strategy, I3-MRec generalizes effectively across various modality-missing scenarios.
@@ 2.3. Information Bottleneck
The Information Bottleneck (IB) principle, based on information theory, is widely applied in machine learning tasks such as model robustness (Wu et al., 2020a; Zhang et al., 2022; Ding et al., 2023), fairness (Gronowski et al., 2023), explainability (Espinosa Zarlenga et al., 2022; Wang et al., 2023c), and recommendation systems (Wei et al., 2022; Yang et al., 2024, 2025; Zhao et al., 2025). For input data $X$, hidden representation $Z$, and prediction label $Y$, IB principle suggests that an effective representation retains minimal sufficient information for the prediction task (Tishby et al., 2000; Tishby and Zaslavsky, 2015; Hu et al., 2024): $Max:I(Y;Z)-\beta I(X;Z)$. where $I(Y;Z)$ and $I(X;Z)$ are the mutual information between variables, with $\beta$ balancing the two terms. Calculating mutual information for continuous variables is difficult, especially in deep learning. Approximations using neural networks, such as MINE (Belghazi et al., 2018), InfoNCE (Oord et al., 2018), and variational methods (Tishby and Zaslavsky, 2015), are commonly used. Recently, CLUB (Cheng et al., 2020) has emerged to estimate the upper bound of mutual information using a log-ratio contrastive loss, which is more general for high-dimensional tasks without prior assumptions.
@@ 3. Preliminaries
Figure 2. Overview of the proposed I3-MRec framework. The model first learns preference-oriented user/item representations using a graph-based approach, guided by IRM to ensure invariance across modalities. Then, an IB-based fusion module generates compact yet effective representations by maximizing preference-relevant information while compressing modality-specific information.
@@ 3.2. Modality Missingness Problem
MRS typically models user preferences by first extracting modality features and then fusing them into a unified item representation. Formally, this process can be described as:
@@ 4.1.1. Collaborative Modality Graph Learning
where $\bm{A}$ is the adjacency matrix built from user-item interactions $\mathcal{R}$ and its transpose $\mathcal{R}^{T}$, and $\bm{D}$ is the corresponding diagonal degree matrix. By aggregating node features across $L$ propagation layers, the final preference-oriented embeddings can be obtained:
@@
Here, $\bm{E}^{k}=\{\bm{E}^{k}_{U},\bm{E}^{k}_{I}\}$ represents preference-oriented representations of users and items. Based on the learned representations, we can estimate the preference score of the user $u$ for item $i$ in the $k$-th modality environment:
@@ 4.1.2. Invariant Learning Optimization
After obtaining predictions across $K$ modality environments, we explicitly encourage the consistency and stability of the learned representations across modalities. Specifically, we employ Invariant Risk Minimization  (Arjovsky et al., 2019; Ahuja et al., 2020; Du et al., 2022) to guide the learning of preference-oriented representations that remain invariant across environments. The corresponding optimization objective is formulated as follows:
@@ 4.2. Missing-aware Feature Fusion
To ensure robust performance under modality-missing scenarios, we design a missing-aware feature fusion module grounded in the Information Bottleneck (IB) principle, as illustrated in the lower part of Figure 3. As discussed above, modality-specific information can become unreliable in such scenarios, which may hinder accurate user preference prediction. To address this, the module is guided to preserve preference-relevant information while compressing modality-specific information, thereby generating compact yet effective multimodal representations.
@@ 4.2.1. Modality-Missing Scenarios Simulation
This strategy exposes the proposed model to various modality-missing scenarios during training, thus encouraging the fusion module to generate robust multimodal representations that ensure stable and consistent recommendation performance across different modality-missing conditions.
@@ 4.2.2. Information Bottleneck-guided Feature Fusion
Given generated representations, we employ the Information Bottleneck principle (Tishby and Zaslavsky, 2015; Tishby et al., 2000; Hu et al., 2024) to guide the optimization of the feature fusion module. It explicitly balances mutual information maximization and minimization, encouraging the model to retain task-relevant signals while compressing modality-specific information. Inspired by previous works (Wei et al., 2022; Liu et al., 2023b; Yang et al., 2024), the Information Bottleneck optimization objective in recommender systems can be formulated as:
@@
Term1: Maximizing Mutual Information with Recommendation Signals. This term aims to maximize the mutual information between generated representations and the recommendation target. Given the user embedding $e_{u}$ and the generated representation $z_{i}$, the predicted preference score is computed as $\bar{r}_{u,i}^{k}=e_{u}^{T},z_{i}$. We select popular BPR ranking loss  (Rendle et al., 2009) to optimize this term:
@@
Term3: Minimizing Mutual Information for Redundancy Reduction. This term aims to compress the negative impact of modality-specific information by minimizing the mutual information between the generated representation and the raw modality input. In other words, it reduces the model’s reliance on potentially unreliable modality information. However, directly computing mutual information between two high-dimensional variables is intractable. To address this, we adopt the Contrastive Log-ratio Upper Bound (CLUB) (Cheng et al., 2020) in our work, which uses a variational distribution $q_{\phi}^{k}(\cdot|\cdot)$ for each modality to approximate the mutual information:
(14) $\mathcal{L}_{comp}=\frac{1}{|K|}\sum_{k\in\{a,t,v\}}{\sum_{i\in\mathcal{I}}\left[\log q_{\phi^{k}}^{k}(z_{i}|e_{i}^{k})-\log q_{\phi^{k}}(z_{i}|x_{i}^{k})\right]}$
here, $q_{\phi}^{k}(\cdot|\cdot)$ is implemented as a two-layer MLP with modality-specific parameter $\phi^{k}$, optimized in a sample-based manner. Minimizing $\mathcal{L}^{comp}$ effectively reduces the task-irrelevant and modality information, yielding a more compact and effective representation.
@@
This unified loss function integrates four components: the recommendation loss ($\mathcal{L}_{rec}$), the invariant representation learning loss ($\mathcal{L}_{IRM}$), the preference consistency loss ($\mathcal{L}_{pcon}$), and the information compression loss ($\mathcal{L}_{comp}$). The hyperparameters $\alpha$ and $\beta$ control the relative importance of the mutual information regularization terms. By jointly optimizing these objectives, the model learns compact and effective preference representations that generalize well to modality-missing scenarios.